,On December 31, 2017, Dow Steel Corporation had 640,000 shares of common stock and 34,000 shares of 8%, noncumulative, nonconvertible preferred stock issued and outstanding. Dow issued a 4% common stock dividend on May 15 and paid cash dividends of $440,000 and $73,000 to common and preferred shareholders, respectively, on December 15, 2018.

On February 28, 2018, Dow sold 57,000 common shares. In keeping with its long-term share repurchase plan, 6,000 shares were retired on July 1. Dow's net income for the year ended December 31, 2018, was $2,300,000. The income tax rate is 40%.
  
Required:
Compute Dow's earnings per share for the year ended December 31, 2018. (Do not round intermediate calculations. Enter your answers in thousands.)

Numerator / Denominator = Earnings per share
/ =
Answer #1
Calculation of total weighted average no of shares
Particulars No of shares
Total no of common stock 665600
add: weigheted avg no of share sold 49400
less: weighted avg no of share retired -3000
Total weighted avg no of shares 712000
Total no of common stock = share held+(no of share held*stock dividend%)
640000 share +640000*4% =665600
weighted avg no of share sold = (period*no of share sold)*stock dividend
57000 share*10/12*1.04 = 49400
weighted avg no of share retired = no of share retired*time
6000share *6/12 = 3000 shares
Earning per share = net income-preferred dividend/weighted avg no of common share outsttanding
earning per share = 2300000-73000/712000
earning per share = 3.127